1. A charge particle with mass 0.05 mg and charge 0.0050 C moves through a uniform magnetic field of strength 1.000 T. If the particle moves at a speed of 90 m/s, what is the maximum value of the magnetic force that can act on the particle?
Answer #1

Charge on the particle is q=0.0050\ C

Magnetic field strength is B=1.000T

Speed of the particle is v=90m/s

Magnetic force on a moving charge is

\vec F=q\vec v\times\vec B

The magnitude of the force is

F=qvB\sin\theta

where \theta is the angle between velocity and the magnetic field.

The maximum value of sin(x) is 1. The force is maximum when the value of sine function is 1.

F_{max}=qvB

F_{max}=(0.0050\ C)(90\ m/s)(1.000\ T)

\boxed{ F_{max}=0.45\ N}